Suppose S and C are subset of R^3. S is the unit open sphere with centre at the origin and C is the open cube= { P(x,y,z)| -1<x<1, -1<y<1, -1<z<1. Which of the following is true. Justify your answer
1. S⊂C 2. C⊂S
1
Expert's answer
2021-07-08T13:57:36-0400

Let us show that "S\subset C." If "P(x,y,z)\in S," then "x^2+y^2+z^2<1." Since "x^2\ge 0" and "y^2+z^2\ge 0," the inequality "x^2+y^2+z^2<1" implies "x^2<1." It follows that "-1<x<1." By analogy, "y^2\ge 0", "x^2+z^2\ge 0" and "x^2+y^2+z^2<1" imply "y^2<1," and hence "-1<y<1." In the same way, "-1<z<1." We conclude that "P(x,y,z)\in C," and hence "S\subset C."